OpenStudy (anonymous):

Sven can type 47 words per minute. How many words would he type in half an hour (30 minutes)?

OpenStudy (anonymous):

Ratio problem. 47 / 1 = x / 30 solve for x.

OpenStudy (anonymous):

47 words a minute. 47 * 30= ?

OpenStudy (anonymous):

You can also use dimensional analysis. \[30 minutes \times \frac{47 words}{1 minute} = 30(47 words)\]

OpenStudy (anonymous):

Notice the "minutes" units cancel when multiplying across, leaving you with words as units.

OpenStudy (anonymous):

thanks to all of you guys! I figured it would be 47*30 it just didn't seem right. so I did it correct thank you and thank you for the unit conversion way as well that's how i originally wanted to start it but for some reason could not figure it out

OpenStudy (anonymous):

:)